# KCL-PH-TH/2026-03

### v2

@@ II The Effective Action of Heterotic String Theory
$\begin{split}\delta S_{0}&=\frac{1}{2\kappa^{2}}\int\left(\frac{1}{2}RT-T^{\mu\nu}R_{\mu\nu}-\frac{\kappa^{2}}{6}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}T+\kappa^{2}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}T^{\mu\nu}+2\kappa^{2}F_{\nu\lambda}\nabla_{\mu}\mathcal{H}^{\mu\nu\lambda}\right)\sqrt{-g}\,d^{D}x\\ &=\int\bigg[\frac{1}{4\kappa^{2}}\left(B_{1}+(D-2)B_{2}\right)R^{2}-\frac{B_{1}}{2\kappa^{2}}R_{\mu\nu}R^{\mu\nu}+\frac{1}{4}\left(B_{3}+(D-2)B_{4}-\frac{B_{1}}{3}-\frac{(D-6)}{3}B_{2}\right)\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\\ &\phantom{=\int}+\frac{1}{2}(B_{1}-B_{3})\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}+\frac{\kappa^{2}}{2}B_{3}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}-\frac{\kappa^{2}}{12}\left(B_{3}+(D-6)B_{4}\right)(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2}\\ &\phantom{=\int}+B_{5}\nabla_{\mu}\mathcal{H}^{\mu\lambda\rho}\nabla^{\nu}\mathcal{H}_{\nu\lambda\rho}\bigg]\sqrt{-g}\,d^{D}x.\end{split}$ (13)
@@
$\begin{split}(\delta S_{0}+S_{1})&=\int\bigg\{\frac{1}{16\kappa^{2}}\left(R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}+\underbrace{\left(\lambda_{1}-8B_{1}\right)}_{\hat{\lambda}_{1}}R_{\mu\nu}R^{\mu\nu}+\underbrace{\left(\lambda_{2}+4B_{1}+4(D-2)B_{2}\right)R^{2}}_{\hat{\lambda}_{2}}\right)+\\ &\phantom{=}-\frac{1}{8}\bigg[\left(1-\frac{1}{3\sqrt{3}}\right)\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\underbrace{\left(\lambda_{3}-4B_{1}+4B_{3}\right)}_{\hat{\lambda}_{3}}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}+\\ &\phantom{=}+\underbrace{\left(\lambda_{4}-2B_{3}-2(D-2)B_{4}+\frac{2}{3}B_{1}+\frac{2(D-6)}{3}B_{2}\right)}_{\hat{\lambda}_{4}}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\bigg]+\\ &\phantom{=}+\frac{\kappa^{2}}{24}\bigg[\left(1-\frac{2}{9\sqrt{3}}\right)\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}+\underbrace{\left(\lambda_{5}+12B_{3}\right)}_{\hat{\lambda}_{5}}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}+\\ &\phantom{=}+\underbrace{\left(\lambda_{6}-2B_{3}-2(D-6)B_{4}\right)}_{\hat{\lambda}_{6}}(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2}\bigg]+\underbrace{\left(\frac{\lambda_{7}}{2}+B_{5}\right)}_{\hat{\lambda}_{7}}\nabla_{\mu}\mathcal{H}^{\mu\lambda\rho}\nabla^{\nu}\mathcal{H}_{\nu\lambda\rho}\bigg\}\sqrt{-g}\,d^{D}x.\end{split}$ (14)
@@
Therefore, the seven coefficients $\lambda_{1},\ldots\lambda_{7}$ are specified within a five parameter space spanned by $B_{1},\ldots,B_{5}$. One particular set of coefficients that yields a unitary action that matches string amplitudes is given by [23]: $\lambda_{1}=-4,\lambda_{2}=1,\lambda_{3}=-2,\lambda_{4}=\frac{1}{3},\lambda_{5}=-3,\lambda_{6}=\frac{2}{3},\lambda_{7}=0$. Plugging these in, we get the most general $\mathcal{O}(\alpha^{\prime})$ effective string action in $D$ spacetime dimensions:
@@
$\begin{split}(\delta S_{0}+S_{1})_{\text{General}}&=\int\bigg\{\frac{1}{16\kappa^{2}}\left(R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-\left(8B_{1}+4\right)R_{\mu\nu}R^{\mu\nu}+\left(1+4B_{1}+4(D-2)B_{2}\right)R^{2}\right)+\\ &\phantom{=}-\frac{1}{8}\bigg[\left(1-\frac{1}{3\sqrt{3}}\right)\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\left(4B_{3}-4B_{1}-2\right)\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}+\\ &\phantom{=\quad\,\,}+\left(\frac{1}{3}+\frac{2}{3}B_{1}+\frac{2(D-6)}{3}B_{2}-2B_{3}-2(D-2)B_{4}\right)\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\bigg]+\\ &\phantom{=}+\frac{\kappa^{2}}{24}\bigg[\left(1-\frac{2}{9\sqrt{3}}\right)\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}+\left(12B_{3}-3\right)\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}+\\ &\phantom{=\qquad}+\left(\frac{2}{3}-2B_{3}-2(D-6)B_{4}\right)(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2}\bigg]+B_{5}\nabla_{\mu}\mathcal{H}^{\mu\lambda\rho}\nabla^{\nu}\mathcal{H}_{\nu\lambda\rho}\bigg\}\sqrt{-g}\,d^{D}x.\end{split}$ (15)
@@
We may restrict these transformations such that the resulting action remains unitary by preserving the Gauss-Bonnet combination $(B_{1}=B_{2}=0)$ and removing the $\nabla_{\mu}\mathcal{H}^{\mu\lambda\rho}\nabla^{\nu}\mathcal{H}_{\nu\lambda\rho}$ term $(B_{5}=0)$. Thus, we obtain:
@@
We will now show (as it has already been done in [30] for example) that the condition of unitarity is incompatible with the Kalb-Ramond field strength having the role of torsion in general $D$ dimensions. To do that, we need to categorize the quadratic curvature invariants in the generalized curvature scheme, i.e. in the case where $H_{\mu\nu\lambda}$ assumes the role of torsion. We thus define a contorted connection by
@@
such that the torsion tensor is given as $T_{\mu\nu\lambda}=\frac{2\kappa}{\sqrt{3}}\mathcal{H}_{\mu\nu\lambda}$. In this scheme, the leading term in the effective action can be written as (see Appendix (A) for more details)
@@
where $\bar{R}$ is the generalized Ricci scalar. In quadratic order, there are $6$ different scalar invariants we may construct, which correspond to the different ways one may contract the generalized Riemann tensor (which is less symmetric than the Levi-Civita connection Riemann tensor), namely:
@@
This “basis” of invariants is not unique, and in fact it is more convenient to work in a different basis, that of [39]:
@@
In this basis, the two Gauss-Bonnet-like terms $\bar{G}^{+}_{4},\bar{G}^{+}_{5}$ appear explicitly. We may calculate all these scalar invariants (which are valid for a connection with general torsion) in our specific case, where the torsion is totally antisymmetric and the Bianchi identity $\nabla_{[\rho}\mathcal{H}_{\mu\nu\lambda]}=0$ also holds. [^5] In our string context, the anomaly terms in the Bianchi identity (7) contribute terms of higher order in $\alpha^{\prime}$ to the effective action, hence they are ignored here. By doing that, we get (modulo total derivatives if we assume that these terms are part of an action):
@@
$\begin{split}\bar{G}^{+}_{2}&=R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}+\frac{2}{3}\kappa^{2}\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}-\frac{4}{3}\kappa^{2}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}+\frac{4}{3}\kappa^{2}(\nabla_{\lambda}H^{\mu\nu\lambda})(\nabla^{\rho}H_{\mu\nu\rho})\\ &-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}+\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma},\end{split}$ (27)
$\begin{split}\bar{G}^{+}_{3}&=R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-2\kappa^{2}\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\frac{4}{3}\kappa^{2}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{4}{3}\kappa^{2}(\nabla_{\lambda}H^{\mu\nu\lambda})(\nabla^{\rho}H_{\mu\nu\rho})\\ &-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}+\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma},\end{split}$ (28)
$\begin{split}\bar{G}^{+}_{4}&=R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-4R_{\mu\nu}R^{\mu\nu}+R^{2}-2\kappa^{2}\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+4\kappa^{2}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{2}{3}\kappa^{2}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\\ &-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}+\frac{1}{9}\kappa^{4}(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2},\end{split}$ (29)
$\begin{split}\bar{G}^{+}_{5}&=R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-4R_{\mu\nu}R^{\mu\nu}+R^{2}+\frac{2}{3}\kappa^{2}\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\frac{4}{3}\kappa^{2}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{2}{3}\kappa^{2}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\\ &-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}+\frac{1}{9}\kappa^{4}(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2},\end{split}$ (30)
@@
$\begin{split}S_{\text{GCS}}&=\int\bigg\{\frac{1}{\kappa^{2}}(A_{4}+A_{5})(R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-4R_{\mu\nu}R^{\mu\nu}+R^{2})+\\ &\phantom{\int}+\left(-2A_{4}+\frac{2}{3}A_{5}\right)\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\left(4A_{4}+\frac{4}{3}A_{5}\right)\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{2}{3}(A_{4}+A_{5})\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\\ &\phantom{\int}+\frac{2}{9}\kappa^{2}\left(32A_{6}-A_{4}-A_{5}\right)\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}-\frac{2}{9}\kappa^{2}\left(16A_{6}+A_{4}+A_{5}\right)\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}+\\ &\phantom{\int}+\frac{1}{9}\kappa^{2}(A_{4}+A_{5})(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2}\bigg\}\sqrt{-g}\,d^{D}x.\end{split}$ (32)

### v3

@@ II The Effective Action of Heterotic String Theory
$\begin{split}\delta S_{0}&=\frac{1}{2\kappa^{2}}\int\bigg(\frac{1}{2}RT-T^{\mu\nu}R_{\mu\nu}-\frac{\kappa^{2}}{6}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}T+\kappa^{2}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}T^{\mu\nu}+\\ &+2\kappa^{2}F_{\nu\lambda}\nabla_{\mu}\mathcal{H}^{\mu\nu\lambda}\bigg)\sqrt{-g}\,d^{D}x\\ &=\int\bigg[\frac{1}{4\kappa^{2}}\left(B_{1}+(D-2)B_{2}\right)R^{2}-\frac{B_{1}}{2\kappa^{2}}R_{\mu\nu}R^{\mu\nu}+\\ &+\frac{1}{4}\left(B_{3}+(D-2)B_{4}-\frac{B_{1}}{3}-\frac{(D-6)}{3}B_{2}\right)\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R+\\ &\phantom{=\int}+\frac{1}{2}(B_{1}-B_{3})\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}+\frac{\kappa^{2}}{2}B_{3}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}-\\ &-\frac{\kappa^{2}}{12}\left(B_{3}+(D-6)B_{4}\right)(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2}+B_{5}\nabla_{\mu}\mathcal{H}^{\mu\lambda\rho}\nabla^{\nu}\mathcal{H}_{\nu\lambda\rho}\bigg]\sqrt{-g}\,d^{D}x.\end{split}$ (13)
@@
$\begin{split}(\delta S_{0}+S_{1})&=\int\bigg\{\frac{1}{16\kappa^{2}}\left(R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}+\underbrace{\left(\lambda_{1}-8B_{1}\right)}_{\hat{\lambda}_{1}}R_{\mu\nu}R^{\mu\nu}+\underbrace{\left(\lambda_{2}+4B_{1}+4(D-2)B_{2}\right)R^{2}}_{\hat{\lambda}_{2}}\right)+\\ &\phantom{=}-\frac{1}{8}\bigg[\left(1-\frac{1}{3\sqrt{3}}\right)\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\underbrace{\left(\lambda_{3}-4B_{1}+4B_{3}\right)}_{\hat{\lambda}_{3}}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}+\\ &\phantom{=}+\underbrace{\left(\lambda_{4}-2B_{3}-2(D-2)B_{4}+\frac{2}{3}B_{1}+\frac{2(D-6)}{3}B_{2}\right)}_{\hat{\lambda}_{4}}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\bigg]+\\ &\phantom{=}+\frac{\kappa^{2}}{24}\bigg[\left(1-\frac{2}{9\sqrt{3}}\right)\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}+\underbrace{\left(\lambda_{5}+12B_{3}\right)}_{\hat{\lambda}_{5}}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}+\\ &\phantom{=}+\underbrace{\left(\lambda_{6}-2B_{3}-2(D-6)B_{4}\right)}_{\hat{\lambda}_{6}}(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2}\bigg]+\\ &+\underbrace{\left(\frac{\lambda_{7}}{2}+B_{5}\right)}_{\hat{\lambda}_{7}}\nabla_{\mu}\mathcal{H}^{\mu\lambda\rho}\nabla^{\nu}\mathcal{H}_{\nu\lambda\rho}\bigg\}\sqrt{-g}\,d^{D}x.\end{split}$ (14)
@@
Therefore, the seven coefficients $\lambda_{1},\ldots,\lambda_{7}$ are specified within a five-parameter space spanned by $B_{1},\ldots,B_{5}$. One particular set of coefficients that yields a unitary action that matches string amplitudes is given by $\lambda_{1}=-4,\lambda_{2}=1,\lambda_{3}=-2,\lambda_{4}=\frac{1}{3},$ $\lambda_{5}=-3,\lambda_{6}=\frac{2}{3},\lambda_{7}=0$ [23]. Plugging these in, we get the most general $\mathcal{O}(\alpha^{\prime})$ effective string action in $D$ spacetime dimensions:
@@
$\begin{split}(\delta S_{0}+S_{1})_{\text{General}}&=\int\bigg\{\frac{1}{16\kappa^{2}}\left(R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-\left(8B_{1}+4\right)R_{\mu\nu}R^{\mu\nu}+\left(1+4B_{1}+4(D-2)B_{2}\right)R^{2}\right)+\\ &\phantom{=}-\frac{1}{8}\bigg[\left(1-\frac{1}{3\sqrt{3}}\right)\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\left(4B_{3}-4B_{1}-2\right)\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}+\\ &\phantom{=\quad\,\,}+\left(\frac{1}{3}+\frac{2}{3}B_{1}+\frac{2(D-6)}{3}B_{2}-2B_{3}-2(D-2)B_{4}\right)\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\bigg]+\\ &\phantom{=}+\frac{\kappa^{2}}{24}\bigg[\left(1-\frac{2}{9\sqrt{3}}\right)\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}+\left(12B_{3}-3\right)\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}+\\ &\phantom{=\qquad}+\left(\frac{2}{3}-2B_{3}-2(D-6)B_{4}\right)(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2}\bigg]+\\ &+B_{5}\nabla_{\mu}\mathcal{H}^{\mu\lambda\rho}\nabla^{\nu}\mathcal{H}_{\nu\lambda\rho}\bigg\}\sqrt{-g}\,d^{D}x.\end{split}$ (15)
@@
We may restrict these transformations such that the resulting action remains unitary by preserving the Gauss–Bonnet combination $(B_{1}=B_{2}=0)$ and removing the $\nabla_{\mu}\mathcal{H}^{\mu\lambda\rho}\nabla^{\nu}\mathcal{H}_{\nu\lambda\rho}$ term $(B_{5}=0)$. Thus, we obtain
@@
We will now show (as has already been accomplished in [30], for example) that the condition of unitarity is incompatible with the Kalb–Ramond field strength, having the role of torsion in general $D$ dimensions. To accomplish that, we need to categorize the quadratic curvature invariants in the generalized curvature scheme, i.e., in the case where $H_{\mu\nu\lambda}$ assumes the role of torsion. We thus define a contorted connection by
@@
such that the torsion tensor is given as $T_{\mu\nu\lambda}=\frac{2\kappa}{\sqrt{3}}\mathcal{H}_{\mu\nu\lambda}$. In this scheme, the leading term in the effective action can be written as follows (see Appendix (A) for more details):
@@
where $\bar{R}$ is the generalized Ricci scalar. In quadratic order, there are six different scalar invariants we may construct, which correspond to the different ways one may contract the generalized Riemann tensor (which is less symmetric than the Levi–Civita connection Riemann tensor), namely as follows:
@@
This “basis” of invariants is not unique, and in fact, it is more convenient to work in a different basis, that of [39], as follows:
@@
In this basis, the two Gauss–Bonnet-like terms $\bar{G}^{+}_{4},\bar{G}^{+}_{5}$ appear explicitly. We may calculate all these scalar invariants (which are valid for a connection with general torsion) in our specific case, where the torsion is totally antisymmetric and the Bianchi identity $\nabla_{[\rho}\mathcal{H}_{\mu\nu\lambda]}=0$ also holds [^5] In our string context, the anomaly terms in the Bianchi identity (7) contribute terms of higher order in $\alpha^{\prime}$ to the effective action; hence, they are ignored here.. By doing that, we obtain modulo total derivatives if we assume that these terms are part of an action:
@@
$\bar{G}^{+}_{2}$ $=R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}+\frac{2}{3}\kappa^{2}\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}-\frac{4}{3}\kappa^{2}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}+\frac{4}{3}\kappa^{2}(\nabla_{\lambda}H^{\mu\nu\lambda})(\nabla^{\rho}H_{\mu\nu\rho})$
$-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}+\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma},$ (27)
$\bar{G}^{+}_{3}$ $=R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-2\kappa^{2}\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\frac{4}{3}\kappa^{2}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{4}{3}\kappa^{2}(\nabla_{\lambda}H^{\mu\nu\lambda})(\nabla^{\rho}H_{\mu\nu\rho})$
$-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}+\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma},$ (28)
$\bar{G}^{+}_{4}$ $=R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-4R_{\mu\nu}R^{\mu\nu}+R^{2}-2\kappa^{2}\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+4\kappa^{2}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}$
$-\frac{2}{3}\kappa^{2}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}$
$+\frac{1}{9}\kappa^{4}(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2},$ (29)
$\bar{G}^{+}_{5}$ $=R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-4R_{\mu\nu}R^{\mu\nu}+R^{2}+\frac{2}{3}\kappa^{2}\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\frac{4}{3}\kappa^{2}\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}$
$-\frac{2}{3}\kappa^{2}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}-\frac{2}{9}\kappa^{4}\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}$
$+\frac{1}{9}\kappa^{4}(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2},$ (30)
@@
$S_{\text{GCS}}$ $=\int\bigg\{\frac{1}{\kappa^{2}}(A_{4}+A_{5})(R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-4R_{\mu\nu}R^{\mu\nu}+R^{2})+$
$\phantom{\int}+\left(-2A_{4}+\frac{2}{3}A_{5}\right)\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\left(4A_{4}+\frac{4}{3}A_{5}\right)\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}$-
$-\frac{2}{3}(A_{4}+A_{5})\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R+\frac{2}{9}\kappa^{2}\left(32A_{6}-A_{4}-A_{5}\right)\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}$-
$-\frac{2}{9}\kappa^{2}\left(16A_{6}+A_{4}+A_{5}\right)\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}+$
$\phantom{\int}+\frac{1}{9}\kappa^{2}(A_{4}+A_{5})(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2}\bigg\}\sqrt{-g}\,d^{D}x.$ (32)